OpenStudy (dls):

Two uniform circular discs having the same mass and same thickness are made from different materials.The disc with smaller Moment of inertia is: A)The one made from more dense material B)The one made from less dense material C)Neither,Both have same MOI D)The disc with larger angular velocity

OpenStudy (dls):

@RnR

OpenStudy (anonymous):

So jaa ab. :/

OpenStudy (dls):

:/ bata dena :| bahut purana doubt ha.

OpenStudy (anonymous):

I dont think it depends on thickness?

OpenStudy (dls):

it does,VERY VERY VERY indirectly.

OpenStudy (anonymous):

Hmm. I dont know.

OpenStudy (dls):

I=MR^2 I=Volume x Density x R^2 I=Area x Thickness x Density x R^2!!!

OpenStudy (anonymous):

Galat hai. Same material? Same area?

OpenStudy (dls):

I=MR^2 M same for both,so only depends on radius?i guess

OpenStudy (anonymous):

Yup. I think so.

OpenStudy (dls):

kiska hoga +?

OpenStudy (anonymous):

Pta nhi yr. Dimaag ki dahi ho chuki hai. :/

OpenStudy (dls):

:/

OpenStudy (anonymous):

@Mashy

OpenStudy (anonymous):

@Mashy Sir?

OpenStudy (dls):

T h i s n i g h t i s s p a r k l i n g Don't you let it go. . I'm wonder stuck .. blushing all the way home... I'll spend f o re v er wondering if you knew.. I was enchanted to meet you oooo..:")

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

wait lemme htnk!

OpenStudy (dls):

htnk

OpenStudy (anonymous):

ok.. since they have to be equally thick and of same mass.. you can imagine one having a large radius (disc a) and one more have small radius (disc b).. definitely the one with larger radius has the higher moment of inertia.. now think which one is more dense :P

OpenStudy (anonymous):

I didnt read it, but there you go with the medal. :P

OpenStudy (dls):

why imagine with different radius?

OpenStudy (anonymous):

lol you do know its very rare that i get it right the first time :P and DLS when you consider a general case you should always consider them having adifferent parameters!

OpenStudy (anonymous):

Will density have any affect at all?

OpenStudy (dls):

Density increases Area Decreases Radius Decreases Inertia decreases

OpenStudy (anonymous):

What do you mean by adjacent parameters?

OpenStudy (dls):

Density increases MOI decreases

OpenStudy (anonymous):

what are you blabbering DLS?? i mean alwys consider the two objects to have different parameters of their characteristics.. since they have same mass and same thickness.. the only thing they can now have different is radius..

OpenStudy (anonymous):

in my earlier post.. disc a is the answer...so now find out whether disc a has bigger density or smaller density!

OpenStudy (dls):

Density increases Area Decreases Radius Decreases Inertia decreases

OpenStudy (dls):

So A)

OpenStudy (anonymous):

IN MY POST DISC A HAS HIGHER RADIUS and Same mass as Disc b which has smaller radius and same mass they have same thicknes.. so which one has more density!?!?? a or b??

OpenStudy (dls):

But why is.. \[\LARGE Density =\frac{k}{Area}?\]

OpenStudy (anonymous):

think logically man.. and tell me which has higher density ? the one with bigger radius or smaller radius? :-/

OpenStudy (dls):

smaller

OpenStudy (anonymous):

obviously right.. the smaller radius has bigger density.. and the bigger radius has smaller density.. and this one which has higher moment of inertia.. so the answer is the one with smaller density!

OpenStudy (dls):